# Addressing Source Scale Bias via Image Warping for Domain Adaptation

### v1

@@ 5 Experiments
5.1 Domain Adaptive Object Detection
Tables 1, 2, 3 and 4 show improvements of our method over state-of-the-art 2PCNet [17] across lighting, weather and geographic adaptations respectively. Moroever, our instance-level saliency guidance always improves over approximate saliency guidance [36, 9]. Note that 2PCNet [17] focused solely on day night adaptation with it’s NightAug augmentation, which was only used in the BDD100K day night experiment (Table 3) and omitted for others. Our method to address scale bias is more general in comparison.
5.2 Domain Adaptive Semantic Segmentation
To demonstrate our method is task agnostic, we incorporate it within domain adaptive semantic segmentation using the popular DAFormer [12] algorithm. The resulting approach shows significant improvement on both Val and Test Splits of the ACDC dataset (Table 5). Here, saliency guidance choice matters: our instance-level image saliency guidance outperforms Static Prior [36] and Geometric Prior [9]. Being agnostic to the DA algorithm, our method also boosts DAFormer extensions like HRDA [13] and MIC [14] (See Supplementary Section B). These extensions mitigate scale bias to some extent but add significant training memory (+5GB) and inference latency (+800 ms on RTX 4090 GPU), whereas our approach requires negligible additional training memory (+0.04 GB) and adds no additional inference latency.
Table 5 also shows our method’s superiority across scenarios, with or without unsupervised domain adaptation. This suggests that the learned backhone features due to our method generalize better to unseen domains. Additionally, Figure 7 illustrates superior performance of our method in identifying both background elements (e.g., sky, sidewalks, buildings) and foreground objects (e.g., traffic lights and traffic signs).
5.3 Visual Diagnosis of Learned Models
In Section 5.2, we noted that the learned backbone feature of our method generalize better. To corroborate this claim, we employed GradCAM [33] visualizations, showcasing ResNet features from detectors trained on BDD100K [47] in Figure 8. The heatmap for models trained with 2PCNet without warping reveals a dispersed focus across the image, indicating a lack of precision. The choice of saliency guidance plays a critical role: warping guided by Geometric Prior [9] results in less focus on the target (e.g., cars) and unnecessary background emphasis. Conversely, our instance-level saliency guidance achieves a focused heatmap on relevant objects, minimizing irrelevant activations.
5.4 Ablation Study
@@
For all ablation studies, we utilize the Cityscape ACDC domain adaptive semantic segmentation task.
Saliency Product and Upperbound . Varying and shows optimal performance at and , according to quantitative scores in Table 7(a). Hence, and are adopted for all major experiments.
Warping on Source vs. Target. Given the necessity of ground truth labels for our instance-level saliency guidance, warping in the target domain for unsupervised domain adaptation is infeasible. Nevertheless, as shown in Table 8 warping source domain images with our saliency guidance surpasses performance using approximated saliency [36, 9] with both source and target domain warping.
Warping with From-Seg vs. Groundtruth bboxes. For object detection, we utilize ground truth bounding boxes for saliency guidance. However, semantic segmentation tasks often lack this data, leading us to use bounding boxes computed from segmentations, “from-seg” for short. Results indicate superior performance with ground truth boxes, but “from-seg” boxes are a close second, highlighting their effectiveness when ground truth is unavailable (Table 8).
Distribution Balancing by Undersampling. We analyzed undersampling strategies that balance the percentages of small, medium, and large objects, aiming to create a more uniform distribution that mitigates scale bias. As shown in Table 7(b), our experiments applying 25% and 50% of undersampling to the Cityscapes dataset indicates that models trained on these “uniform” subsets perform worse than those trained on the complete dataset, highlighting the challenges of addressing scale bias through simple subsampling.
@@ 6 Limitations and Conclusion
In this work, we address scale bias with attentional processing, using saliency-guided image warping that integrates seamlessly into any unsupervised domain adaptation method. Leveraging scale statistics from the source dataset and ground truth labels, our approach effectively navigates challenges in lighting, weather, and geographic differences. Extensive experiments demonstrate the effectiveness of our warping method and instance-level saliency guidance. Importantly, our approach adds no inference latency. Depending on the available information (e.g., vanishing point, target statistics), other priors could be used with our approach, showing improvement over baselines, albeit to a lesser extent.
One limitation of our method arises in densely populated scenes (e.g., New York’s Times Square), where the limited background space challenges the expansion and contraction capabilities of our warping. Besides, our approach is less effective on certain synthetic datasets, such as Synthia [29], which do not exhibit realistic scale bias. However, our approach proves effective on other synthetic datasets like GTA5 [28], a role-playing game dataset featuring city driving components with realistic traffic and exhibiting realistic scale biases.
Societal Impact. While our method demonstrates improved perception on publicly available night and adverse weather driving datasets, extensive testing is required before real-world deployment on self-driving vehicles.
Acknowledgment. This work was supported in parts by a sponsored research contract from General Motors Research-Israel, NSF grant CNS-2038612, and DOT RITA Mobility-21 Grant 69A3551747111.
@@ Appendix A Synthetic to Real Domain Adaptation
Our proposed method addresses source scale biases that are present in natural scene datasets. Synthetic To Real (Sim2Real) Domain Gap is not a focus of our work as a synthetic scene dataset may or may not exhibit the same bias.
However, we do explore if our method improves performance on the Sim2Real Domain Gap by considering GTA [28] and Synthia [29] synthetic datasets. Table 9 and Table 10 show the results for GTA Cityscapes and Synthia Cityscapes domain adaptation respectively.
Our method improves GTA Cityscapes adaptation considerably, while improvement on Synthia Cityscapes adaptation is marginal, which corresponds to our intuition regarding both the source datasets (See Fig 9).
Grand Theft Auto 5 (GTA) is a role playing game with a city driving component along with realistic traffic, thus GTA images are close to real driving scenes and hence exhibits similar scale biases. Synthia, on the other hand, is a rendered dataset of a virtual city, and does not have realistic traffic simulation of dynamic objects like cars and people. Morever, many images are not captured from a vehicle’s perspective. Thus, images are not close to real driving scenes and do not exhibit scale bias.
@@ Appendix B Comparison with other Domain Adaptation Strategies
Domain adaptation extensions have been proposed on the DAFormer [12] training framework, such as HRDA [13] and MIC [14] which also show good performance.
@@
Differences from DAFormer. HRDA [13] proposed a multi-scale high-resolution crop training strategy along with sliding window inference. MIC [14] proposed a masking consistency strategy for target domain images to improve spatial context.
HRDA [13] departs from DAFormer [12] by increasing training and inference image scales by a factor of 2 and performing sliding window inference (a test time scale augmentation). Training and inference computational costs significantly increase with HRDA’s strategy [13]. Correspondingly, our inference costs are same as DAFormer [12] (as we do not warp during test-time) while training time and memory consumption is marginally higher (See Table 11). Lastly, instance-level image warping is orthogonal to the aforementioned training strategies and can be used simultaneously with other training strategies.
@@
Evaluation Methodology. Based on the above discussion, we apply HRDA’s and MIC’s strategy on DAFormer, but for a fair comparison, we perform training and inference following DAFormer’s [12] training/testing image scales and evaluation paradigm. The alternative comparison would be to train and evaluate on full-scale images for parity with HRDA, however, we do not have a GPU with high memory availability (like NVIDIA Tesla A40 with 48 GB or Tesla A100 with 80 GB memory).
@@
Results. Table 12 and 13 present results for Cityscapes DarkZurich and Cityscapes ACDC respectively. MIC (HRDA) corresponds to MIC and HRDA training strategies added to DAFormer. We compare with DAFormer and MIC (HRDA) with and without including our Instance-level image warping approach. We observe smaller improvements in conjunction with our method with MIC (HRDA) as HRDA uses multi-scale cropping strategies, which also attempts to alleviate scale bias to some extent.
As our instance-level image warping approach is orthogonal to other training strategies and can be used in plug-and-play manner, we studied the interaction of our method with MIC [14] and HRDA [13] individually (See Table 14).
We observed that our method when employed in conjunction with MIC and HRDA consistently improved performance, however, was not able to improve upon the scenario where our method is applied to DAFormer directly. Moreover, HRDA failed to improve upon DAFormer itself when trained on DAFormer’s training scale (1024 512) instead of full scale (2048 1024).
We concluded that high resolution training is necessary for HRDA as posited [13] and the strategy does not work at all image scales. Further, we observed HRDA performs two kinds of cropping – a HR-detail crop to better segment small object details and LR-context crop to adapt better to large objects. At DAFormer’s training scales, the cropped size of the detail crop would be very small and hurt performance. The advantage of our instance-level warping method is that it increases focus on small object details, thus we replaced detail crop with our instance level image warping. Removing detail crop from HRDA improved performance over DAFormer itself, and in conjunction with our method showed the best performance.
@@
HRDA (DAFormer) | 58.3 | 68.5 | 59.4 | 82.8 | 50.4 | 40.8 | 50.3 | 42.4 | 44.4 | 77.6 | 38.0 | 69.4 | 55.7 | 27.2 | 83.2 | 77.4 | 78.2 | 79.0 | 34.9 | 48.4 | ||||||||||
HRDA (DAFormer) + Ours | 62.1 | 89.7 | 61.1 | 83.9 | 43.4 | 39.5 | 52.7 | 43.1 | 45.0 | 75.6 | 38.7 | 86.1 | 55.0 | 28.0 | 84.9 | 81.1 | 88.5 | 86.0 | 44.9 | 53.4 |
@@ Appendix C Supervised Setting
In Section 3.2, we mentioned that source pre-training also improves performance in the supervised setting on the source domain itself. We show improvements in segmentation performance on Cityscapes, when applied to SegFormer model (base architecture of the DAFormer [12] training strategy), see Table 15. Please observe visual results of our method in Figure 10. We also show improvements in detection performance on BDD100K (Day) and BDD100 (Clear) when applied to Faster R-CNN (base architecture of 2PCNet [17] training strategy), see Table 16 and Table 17. BDD100K (Day) includes both clear and bad weather images taken during the day, while BDD100K (Clear) includes both day and night images taken in clear weather.
@@ Appendix D Additional Domain Adaptations
Cityscapes Foggy Cityscapes Table 18 shows the result for domain adaptative object detection from Cityscapes Foggy Cityscapes.
@@ Appendix E Additional Analysis
Pixel Accuracy Visualizations Figure 11 show per-pixel accuracy visualization. For both Cityscapes [4] and ACDC [32], a noticeable predominance of red over blue is observed, showing a clear advantage of our method upon DAFormer [12] for semantic segmentation.
@@
BDD100K Clear Rainy Object Detection. Qualitative comparisons are shown in Figure 12. Our method demonstrates superior object detection under rainy conditions by accurately identifying vehicles and minimizing false positives, such as the misidentified pedestrians and cars evident in the 2PCNet [17] predictions.
@@
Cityscapes DarkZurich Semantic Segmentation. Qualitative comparisons are shown in Figure 13. The proposed method exhibits a segmentation output that is more closely aligned with groundtruth, particularly while predicting road boundaries and consistently identification of urban elements such as sidewalk, terrain, and traffic signs. This indicates that our method has superior domain adaptation capabilities in challenging low-light conditions.
Appendix F Additional Technical Details

### v2

@@ 3.3 Warping for Unsupervised Domain Adaptation
4 Experimental Methodology
We implement our saliency warping technique alongside a base domain adaptation algorithm, ensuring fair comparisons by maintaining the same model, datasets, training schedules, hyperparameters, and random seeds.
Object Detection: Following the recent and popular domain adaptation strategy 2PCNet [17], we use Faster R-CNN [25] with ResNet-50 [11], adhering to their training hyperparameters and protocols. While 2PCNet [17] focuses solely on Day-to-Night adaptation, our method addresses various adaptation scenarios. For scenarios other than Day-to-Night adaptation, we exclude the NightAug augmentation proposed by 2CPNet for both their method and ours.
Semantic Segmentation: We follow DAFormer [13] by employing the same SegFormer [45] head and MiT-B5 [45] backbone, adhereing to their training hyperparameters, protocols, and seed for fair comparison. While Sim2Real Gap result presented in DAFormer [13] are not our focus, relevant results and discussion can be found in Appendix A.
5 Experiments
@@
5.1 Domain Adaptive Object Detection
Tables 1, 2, 3 and 4 show improvements over the state-of-the-art 2PCNet [17] across lighting, weather and geographic adaptations. Moreover, our instance-level saliency guidance consistently outperforms approximate saliency guidance [35, 9]. Note that 2PCNet [17] focused solely on Day Night adaptation with its NightAug augmentation, which was only used in the BDD100K Day Night experiment (Table 3) and omitted for others.
Figure 5 shows examples where 2PCNet [17] mistakes parts of the road for cars and street lights for traffic lights (For other domain adaptations, see Appendix E). In contrast, our method accurately distinguish these foreground objects and background element in challenging lighting and weather conditions.
5.2 Domain Adaptive Semantic Segmentation
To demonstrate the task-agnostic nature of our method, we integrate it into domain adaptive semantic segmentation using the popular DAFormer [13] algorithm. We observe significant improvement on both Val and Test Splits of the ACDC dataset (Table 5). Our instance-level image saliency guidance outperforms Static Prior [35] and Geometric Prior [9]. Being agnostic to the DA algorithm, our method also boosts DAFormer extensions like HRDA [14] and MIC [15] (See Appendix B). These extensions add significant training memory (+5GB) and inference latency (+800 ms on RTX 4090 GPU), whereas our approach requires negligible additional training memory (+0.04 GB) and adds no additional inference latency.
Table 5 also demonstrates our method’s superiority across scenarios, both with and without unsupervised domain adaptation, suggesting the learnt backbone features generalize better to unseen domains. While we briefly discuss some qualitative results using GradCam [31] visualizations, detailed visual diagnosis of the learned backbones is available in Appendix E. Additionally, Figure 6 illustrates the superior performance of our method in identifying both background (e.g., sky, sidewalks, buildings) and objects (e.g., traffic lights, traffic signs).
5.3 Ablation Study
For all the ablations below, we use Cityscape ACDC adaptation, except for ‘Distribution Balancing by Undersampling’, where we fully supervised on Cityscapes.
Saliency Product and Upperbound . Varying and shows optimal performance at and , according to quantitative scores in Table 7(a). Hence, and are used for all major experiments.
Distribution Balancing by Undersampling. We analyze undersampling strategies that select images to create a more uniform object-background distribution. Table 7(b) shows models trained with 25% and 50% undersampling on the Cityscapes dataset perform worse than those trained on the complete dataset, highlighting the ineffectiveness of simple undersampling to address object-background imbalance.
Warping on Source vs. Target. Warping in the target domain for unsupervised domain adaptation is infeasible for instance-level saliency guidance because it requires groundtruth labels. However, Table 8 shows that warping source domain images using our instance-level saliency guidance outperforms using approximated saliency [35, 9] with both source and target domain warping. We also experimented with warping during testing, but this approach introduced additional latency and failed to improve performance compared to warping only the source.
Warping with From-Seg vs. Groundtruth bboxes. For object detection, we utilize ground truth bounding boxes for saliency guidance. However, semantic segmentation tasks often lack this data, leading us to use bounding boxes computed from segmentations, ‘from-seg’ for short. Results indicate superior performance with ground truth boxes, but ‘from-seg’ boxes are a close second, highlighting their effectiveness when ground truth is unavailable (Table 8).
@@ 6 Limitations and Conclusion
In this work, we demonstrate that saliency guided in-place image warping improves Unsupervised Domain Adaptation (UDA). Unlike previous approaches that indiscriminately process objects and backgrounds, our instance-level saliency guidance oversamples object regions and undersamples background regions. This improves learned backbone features, which in turn helps domain adaptation. Extensive experiments validate the effectiveness of our warping approach and instance-level saliency guidance. Notably, our approach is efficient, with minimal training overhead and no additional inference latency, and is agnostic to the adaptation task, target domain, saliency guidance, model architecture, and domain adaptation algorithm.
One limitation of our method is its reduced effectiveness in densely populated scenes (e.g., Times Square) where there is limited background to reduce for enlarging foreground objects. Our in-place warping strategy is less effective when background is less than one-fourth of the image. Besides, our approach is less effective on synthetic datasets like Synthia [27], captured from various viewpoints. However, it proves effective on other synthetic datasets like GTA5 [26], which feature realistic city driving scenes.
@@ Appendix A Synthetic to Real Domain Adaptation
The Synthetic-to-Real (Sim2Real) Domain Gap is not the focus of this work, as synthetic scene datasets may have suboptimal simulations and inconsistent viewpoints, resulting in unrealistic object-background distributions with non-dominant backgrounds, where our strategy of oversampling objects is less effective.
However, we do explore how our method impacts performance on the Sim2Real domain gap by considering the GTA [26] and Synthia [27] synthetic datasets. We expect our method to perform better with realistic datasets and less effectively with unrealistic datasets.
As shown in Fig. 7, Grand Theft Auto V (GTA) is a role-playing game with a city driving component and realistic traffic, making images closely resemble real driving scenes. In contrast, Synthia is a rendered dataset of a virtual city with unrealistic traffic simulation of dynamic objects like cars and people. Additionally, many Synthia images are not captured from a vehicle’s perspective, making them less representative of real driving scenes.
Table 9 shows that our method significantly improves domain adaptation from GTA Cityscapes, while Table 10 shows only marginal improvement for Synthia Cityscapes. This result aligns with our expectations for both source datasets.
@@ Appendix B Comparison with other Domain Adaptation Strategies
B.1 Differences between HRDA/MIC and DAFormer
@@
Domain adaptation extensions have been proposed for the DAFormer [13] training framework, such as HRDA [14] and MIC [15], both of which demonstrate strong performance. HRDA [14] introduces a multi-scale high-resolution crop training strategy combined with sliding window inference, whereas MIC [15] implements a masking consistency strategy for target domain images to enhance spatial context.
B.2 Computational Efficiency
Table 11 shows the computational efficiency comparison using the DAFormer image scale. We observe that incorporating HRDA into DAFormer significantly increases training and inference computational costs due to HRDA’s train-time HR-cropping and test-time sliding window inference. In contrast, our inference costs remain the same as those of DAFormer [13], as we do not warp during test time. Our training memory consumption are only slightly higher, due to the lightweight design of our warping-related modules. Additionally, our saliency-guided image warping is orthogonal to the aforementioned training strategies and can be utilized together with them.
B.3 Evaluation Methodology
@@
Based on the above discussion, we apply HRDA’s and MIC’s strategies to DAFormer. However, for a fair comparison, we perform training and inference following DAFormer’s [13] training/testing image scales and evaluation paradigm. An alternative comparison would involve training and evaluating on full-scale images to match HRDA, which deviates from original DAFormer [13] training and evaluation that uses half-scaled images. Unfortunately, we lack access to a GPU with high memory capacity, such as the NVIDIA Tesla A40 with 48 GB or Tesla A100 with 80 GB of memory.
B.4 Results and Analysis
@@
Cityscapes DarkZurich/ACDC Semantic Segmentation: Table 12 and 13 present results for Cityscapes DarkZurich and Cityscapes ACDC semantic segmentation, respectively. MIC (HRDA) refers to MIC and HRDA training strategies added to DAFormer. We compare DAFormer and MIC (HRDA) with and without our instance-level saliency guided image warping. We observe smaller improvements with our method when combined with MIC (HRDA) due to HRDA’s use of multi-scale cropping. See next paragraph for explanations.
Ablation on Cityscapes ACDC Semantic Segmentation: We studied the interaction of our saliency-guided image warping with MIC [15] and HRDA [14] individually, as our method is orthogonal and plug-and-play. Table 14 shows that (a) HRDA did not improve DAFormer when trained at DAFormer’s scale (1024 512) instead of full scale (2048 1024). (b) Our method improve performance for MIC and HRDA, but neither combination outperform DAFormer+Ours. We investigate and find that our limited improvement on HRDA is due to their use of ‘detail crop’. The HRDA detail crop focuses on small object regions, which our warping already oversamples. This results in redundant efforts and reduce overall effectiveness. To verify this, we remove the detail crop from HRDA and observe that our method showed better improvement on this HRDA∗ variant (see Table 14).
@@
HRDA∗ (DAFormer) | 58.3 | 68.5 | 59.4 | 82.8 | 50.4 | 40.8 | 50.3 | 42.4 | 44.4 | 77.6 | 38.0 | 69.4 | 55.7 | 27.2 | 83.2 | 77.4 | 78.2 | 79.0 | 34.9 | 48.4 | ||||||||||
HRDA∗ (DAFormer) + Ours | 62.1 | 89.7 | 61.1 | 83.9 | 43.4 | 39.5 | 52.7 | 43.1 | 45.0 | 75.6 | 38.7 | 86.1 | 55.0 | 28.0 | 84.9 | 81.1 | 88.5 | 86.0 | 44.9 | 53.4 |
@@ Appendix C Supervised Setting
Supervised setting is not the focus of this work. Instance-level saliency guidance is effective in domain adaptation because its focus on objects significantly reduces the negative impact of backgrounds with large cross-domain variations. In supervised settings, where background variations are low within a single-domain, this focus on objects does not provide the same benefit.
In Section 3.2 of the main paper, we mentioned that source pre-training improves performance in the supervised setting on the source domain. Results are presented below.
Cityscapes Semantic Segmentation: We demonstrate improved semantic segmentation performance on Cityscapes by applying our method to the SegFormer model, which serves as the base architecture for the DAFormer [13] training strategy (see Table 15). Visual comparisons of our method versus SegFormer can be observed in Figure 8.
BDD100K Object Detection: We demonstrate improved object detection performance on BDD100K (Day) and BDD100 (Clear) when applying our method to Faster R-CNN, the base architecture of the 2PCNet [17] training strategy (see Table 16 and Table 17). Note that BDD100K (Day) includes images taken during the day in both clear and bad weather, while BDD100K (Clear) includes images taken in clear weather during both day and night.
@@ Appendix D Additional Domain Adaptations
Cityscapes Foggy Cityscapes: Table 18 shows the result for domain adaptative object detection from Cityscapes Foggy Cityscapes. Our improvement is small on the Foggy Cityscapes dataset. This is because (a) the baseline is already strong when dealing with easy synthetic fog (b) there is little cross-domain background variation, leaving minimal room for improvement with our warping approach that oversamples objects to reduce the negative impacts of cross-domain large background variations.
@@ Appendix E Additional Analysis
Grad-CAM Visualizations of the Learned Model: In Section 5.2 of the main paper, we claimed that the learned backbone features obtained through our training method generalize better. To support this claim, we use GradCAM [31] visualizations to showcase ResNet features from detectors trained on BDD100K [48], as shown in Figure 9. We observe that (a) Heatmaps for models trained with 2PCNet show that with our warping, there is a strong focus on salient objects with minimal distracting activation, while without our warping, the focus is dispersed across the image, indicating a lack of precision. (b) The choice of saliency guidance matters: our instance-level saliency guidance ensures a strong focus on salient objects, whereas warping guided by Geometric Prior [9] results in less emphasis on target objects (e.g., cars) and an unnecessary focus on the background.
Grad-CAM Visualizations for Multiple Objects: Using the same backbone layers, Figure 10 shows that Grad-CAM contributions from background pixels are smaller compared to those from foreground object pixels when our method is used during training. This demonstrates improved focus on foreground objects over the background when trained with our instance-level saliency guidance.
@@
Per-Pixel Accuracy Difference Visualizations: Figure 11 shows per-pixel accuracy difference visualizations. For both Cityscapes [5] and ACDC [30], a noticeable predominance of red over blue is observed, indicating a clear advantage of our method over DAFormer [13] for semantic segmentation.
@@
BDD100K Clear DENSE Foggy Object Detection: Qualitative comparisons are shown in Figure 12. Our method demonstrates superior object detection under real foggy conditions by accurately identifying objects like streetlights and vehicles. In contrast, 2PCNet [17] misidentifies windows as pedestrians, a mistake our approach avoids.
BDD100K (Clear Rainy) Object Detection: Qualitative comparisons are shown in Figure 13. Our method demonstrates superior object detection under rainy conditions by accurately identifying vehicles and minimizing false positives, such as misidentified pedestrians and cars in the 2PCNet [17] predictions.
@@
Cityscapes DarkZurich Semantic Segmentation: Qualitative comparisons are shown in Figure 14. The proposed method produces segmentation outputs that more closely align with the ground truth, particularly in predicting road boundaries and consistently identifying urban elements such as sidewalks, terrain, and traffic signs. This indicates that our method has superior domain adaptation capabilities in challenging low-light conditions.
Appendix F Additional Experimental Details
The following are the datasets used in our experiments.
BDD100K [48] features 100,000 images with a resolution of 1280x720 for object detection and segmentation, covering various weather conditions and times of day, and includes annotations for 10 categories.
Cityscapes [5] provides 5,000 images of urban road scenes at a resolution of 2048x1024 in clear weather for semantic segmentation, with annotations for 19 categories.
ACDC [30] is designed for adverse conditions such as fog and snow, including 1,600 images at a resolution of 2048x1024 for segmentation across 19 categories.
Dark Zurich [28] is tailored for low-light conditions, offering 2,416 unlabeled nighttime images and 151 labeled twilight images for segmentation, all at a resolution of 1920x1080, with a focus on urban settings.
Appendix G Additional Technical Details